OpenStudy (anonymous):

i will give medal The vertices of a triangle are A(16, 0), B(9, 2), and C(0, 0). Using fractions, indicate that the slopes of its sides are: Slope of AB = a0 Slope of BC = a1 Slope of AC = a2 The slopes of its altitudes are: Altitude AB = a3 Altitude BC = a4 Altitude AC = a5 i need the formula for side slopes and altitudes slopes.

OpenStudy (anonymous):

its too long,i'll solve a part,try to solve the rest on your own slope of AB=(yB-yA/xB-xA)=(2-0/9-16)=-2/7 altitude relative to AB is perpendicular to AB so its slope is 7/2

OpenStudy (anonymous):

how did you find its altitude

OpenStudy (anonymous):

how do you find the altitude

OpenStudy (anonymous):

i will give another M3d41

hartnn (hartnn):

product of slope of perpendicular lines =-1 thats why slope of altitude of AB is 7/2...(which is -1/(-2/7))

OpenStudy (anonymous):

-1 divided by -2/7 ???

hartnn (hartnn):

\[\frac{-1}{\frac{-2}{7}}=\frac{7}{2}\]

hartnn (hartnn):

isn't it?

OpenStudy (anonymous):

on the calculator it shows as 3.5

hartnn (hartnn):

7/2 is indeed 3.5

OpenStudy (anonymous):

yes your right one more thing

OpenStudy (anonymous):

is this slope correct? bc= -2/-9 ac - 0/-16

hartnn (hartnn):

-2/-9 is same as 2/9 = 0.2222 .....keep it as 2/9 -0/-16 = 0

OpenStudy (anonymous):

ok thanks so much

hartnn (hartnn):

and yes,u applied the formula correwctly

OpenStudy (anonymous):

yesssssss

hartnn (hartnn):

if u got it tell me the slope of altitude of AC

OpenStudy (anonymous):

you said 0

hartnn (hartnn):

that was slope of AC now i am asking slope of altitude of AC

OpenStudy (anonymous):

0

hartnn (hartnn):

nopes ,remember product of slopes is -1,for perpendicular lines.

OpenStudy (anonymous):

-1/0 ?

hartnn (hartnn):

yesss,-1/0 so what's exatly -1/0 ??

OpenStudy (anonymous):

infinity

OpenStudy (anonymous):

?

OpenStudy (anonymous):

thats what google calculator said

hartnn (hartnn):

infinity is correct :)

OpenStudy (anonymous):

thanks

hartnn (hartnn):

so now did u get all the answers?

OpenStudy (anonymous):

yes but i had to type undefined instead of infinity but everything else was correct